OpenStudy (anonymous):

write an equation for a linear function f that has the given values f(-2)=15, f(1)=9?

OpenStudy (anonymous):

find the slope m = (y2 - y1)/(x2 - x1) = (15-9)/(-2 - 1) = -2 Using point slope we have y - 9 = -2(x - 1)